# Multi-Head Explainer: A General Framework to Improve Explainability in CNNs and Transformers

### v1

@@ Abstract
In this study, we introduce the Multi-Head Explainer (MHEX), a versatile and modular framework that enhances both the explainability and accuracy of Convolutional Neural Networks (CNNs) and Transformer-based models. MHEX consists of three core components: an Attention Gate that dynamically highlights task-relevant features, Deep Supervision that guides early layers to capture fine-grained details pertinent to the target class, and an Equivalent Matrix that unifies refined local and global representations to generate comprehensive saliency maps. Our approach demonstrates superior compatibility, enabling effortless integration into existing residual networks like ResNet and Transformer architectures such as BERT with minimal modifications. Extensive experiments on benchmark datasets in medical imaging and text classification show that MHEX not only improves classification accuracy but also produces highly interpretable and detailed saliency scores.
@@ 4.3 Saliency Score Analysis
Image Classification: MHEX-Net consistently demonstrated superior coverage, closely resembling semantic segmentation, while other methods exhibited characteristic strengths and limitations, such as focusing on localized high-level features or producing fragmented scores.The results are illustrated in Figure 4 (colorectal adenocarcinoma epithelium) and Figure 5 (smooth muscle tissue).
@@
Original MHEX Grad-CAM SHAP Layer-CAM
Figure 4: Saliency score comparison for colorectal adenocarcinoma epithelium. Colorectal adenocarcinoma epithelium is characterized by irregularly arranged epithelial cells with enlarged nuclei and poor glandular differentiation. MHEX provides comprehensive coverage of cancerous regions, effectively capturing diagnostic features.
@@
Original MHEX Grad-CAM SHAP Layer-CAM
Figure 5: Saliency score comparison for smooth muscle tissue. Smooth muscle tissue consists of elongated fibers arranged in parallel, with nuclei distributed along the fibers. MHEX effectively highlights the fibers with uniform coverage, closely resembling semantic segmentation.
@@ 5 Conclusion
Extensive experiments on medical imaging and text classification datasets demonstrate that MHEX enhances both classification performance and saliency map clarity compared to existing methods.
For those who wish to explore the implementation of MHEX, we kindly refer you to the guidelines in Appendix E.
@@ Appendix C More Results
Limpkin
Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption
Sandpiper

### v2

# Multi-Head Explainer: A General Scalable Framework to Improve Explainability in CNNs and Transformers

@@ Abstract
In this study, we introduce the Multi-Head Explainer (MHEX), a versatile and modular framework that enhances both the explainability and accuracy of Convolutional Neural Networks (CNNs) and Transformer-based models. MHEX consists of three core components: an Attention Gate that dynamically highlights task-relevant features, Deep Supervision that guides early layers to capture fine-grained details pertinent to the target class, and an Equivalent Matrix that unifies refined local and global representations to generate comprehensive saliency maps. Our approach demonstrates superior compatibility, enabling effortless integration into existing residual networks like ResNet and Transformer architectures such as BERT with minimal modifications. Extensive experiments on benchmark datasets in medical imaging and text classification show that MHEX not only improves classification accuracy but also produces highly interpretable and detailed saliency scores. [^1] For those who wish to explore the implementation of MHEX, we kindly refer you to the guidelines in Appendix E
@@ 4.3 Saliency Score Analysis
Image Classification: MHEX-Net consistently demonstrated superior coverage, closely resembling semantic segmentation, while other methods exhibited characteristic strengths and limitations, such as focusing on localized high-level features or producing fragmented scores.The results are illustrated in Figure 4 (colorectal adenocarcinoma epithelium and smooth muscle tissue) and Figure 5 (ImageNet).
Original MHEX Grad-CAM SHAP Layer-CAM
@@
Figure 4: Saliency score comparison for smooth muscle tissue and colorectal adenocarcinoma epithelium. Colorectal adenocarcinoma epithelium is characterized by irregularly arranged epithelial cells with enlarged nuclei and poor glandular differentiation. Smooth muscle tissue consists of elongated fibers arranged in parallel, with nuclei distributed along the fibers. MHEX effectively highlights the fibers and cancerous regions with uniform coverage, closely resembling semantic segmentation.
Original Image MHEX Grad-CAM Original Image MHEX Grad-CAM
Limpkin
Refer to caption Refer to caption Refer to caption
Sandpiper
Refer to caption Refer to caption Refer to caption
@@ Appendix E Guidelines
  • Application to Other Networks: We believe that MHEX is applicable to a variety of neural network architectures beyond CNNs and Transformers. For instance, in Graph Neural Networks (GNNs), MHEX should function effectively by treating graph nodes analogously to tokens in transformer-based models. However, it is important to note that standard GNN architectures, such as Graph Convolutional Networks (GCNs) (Kipf & Welling, 2016) and Graph Attention Networks (GATs) (Veličković et al., 2017), typically lack residual connections. Therefore, a key consideration is how to introduce residual links within these frameworks to seamlessly integrate MHEX. Additionally, MHEX can also be applied to medical image segmentation tasks, particularly when used with powerful architectures like UNet and UNet-Transformer (Ronneberger et al., 2015; Petit et al., 2021). Given the different nature of segmentation tasks, the primary modification involves replacing the equivalent matrix ($W_{\text{equiv}}$) with an equivalent convolution ($\text{Conv}_{\text{equiv}}$), allowing MHEX to better adapt to the spatial features inherent in these tasks.